10. (1 point) Use Part I of the Fundamental Theorem of Cal- culus to find the derivative of 3x u+1 g(x) =
Answer #1

g(x)=\int _{9x}^{3x}\frac{u+1}{u-4}du

\Rightarrow g(x)=\int _{9x}^{3x}(\frac{5}{u-4} +1)du

\Rightarrow g(x)=\int _{9x}^{3x}\frac{5}{u-4} du +\int _{9x}^{3x}du

\Rightarrow g(x)=5\int _{9x}^{3x}\frac{du}{u-4} +\int _{9x}^{3x}du

\Rightarrow g(x)=[5\ln \left|u-4\right|+u+C]_{9x}^{3x}

\Rightarrow g(x)=3x+\frac{5}{2}\ln \left(\left(4-3x\right)^2\right)-\left(9x+\frac{5}{2}\ln \left(\left(4-9x\right)^2\right)\right)

\Rightarrow g(x)=-6x+5\ln \left(-3x+4\right)-5\ln \left(-9x+4\right)

therefore, g'(x) will be

g'(x)=\frac{\mathrm{d} }{\mathrm{d} x}[-6x+5\ln \left(-3x+4\right)-5\ln \left(-9x+4\right)]

\Rightarrow g'(x)=-6[\frac{\mathrm{d} }{\mathrm{d} x}x]+5[\frac{\mathrm{d} }{\mathrm{d} x}\ln \left(4-3x\right)]-5[\frac{\mathrm{d} }{\mathrm{d} x}\ln \left(-9x+4\right)]

\Rightarrow g'(x)=-6[\frac{\mathrm{d} }{\mathrm{d} x}x]+5[\frac{1}{4-3x}][\frac{\mathrm{d} }{\mathrm{d} x} \left(4-3x\right)]-5[\frac{1}{4-9x}][\frac{\mathrm{d} }{\mathrm{d} x} \left(-9x+4\right)]

\Rightarrow g'(x)=-6-\frac{15}{4-3x}+\frac{45}{4-9x}